OpenStudy (anonymous):

Find the first 3 nonzero terms in each of two solutions (not multiples of each other) about x = 0

OpenStudy (anonymous):

\[xy'' + y' - y = 0\] I have a first solution no problem, but this DE gives a double root, so I have no idea how to obtain the terms of a second solution. The first several terms of the solution I did find are: \[y_{1}(x) = 1+x+\frac{1}{4}x^{2}+\frac{1}{36}x^{3}+ ...\] The answer for the 2nd solution, which I cannot find for the life of me, is: \[y_{2}(x) = y_{1}(x)lnx - 2x-\frac{3}{4}x^{2}-\frac{11}{108}x^{3}+ ...\] the y2 = y1(x)lnx I know is just a part of the form of a 2nd solution, but those series terms I have no idea. Can anyone show me how to obtain a second series solution and, consequently, those terms that the book has as a solution? x_x

OpenStudy (loser66):

@ikram002p I have a question on x =0. When x =0, P(x) =0 which makes x =0 is not ordinary point, right? so that how to solve it then?

OpenStudy (anonymous):

Its not an ordinary point, correct, its a regular singular point.

OpenStudy (loser66):

No, it not to me :( to check whether it is aregular singular point, I consider: \[lim_{x\rightarrow 0} x*\dfrac{1}{x}=lim_{x\rightarrow 0}1=1\\lim_{x\rightarrow 0}x^2\dfrac{1}{x}=lim_{x\rightarrow 0}x=0\] which give me x =0 is not a regular singular point, :(

OpenStudy (loser66):

@Kainui

OpenStudy (anonymous):

That's fine, as long as the limits are finite, which they are.

OpenStudy (loser66):

I know. but the characteristic equation from those limits is r^2 +(1-1)r + 0 =0 which give us r =0 , and it's nonsense, right?

OpenStudy (anonymous):

That's what its supposed to be o.o You get a double root of r = 0. The one solution I have was derived using r = 0, but theres a way to find a second solution when you have double roots, but Im not sure how to do that. Everything Ive tried has come out wrong.

OpenStudy (loser66):

Ok, ignore it. if I continue your work, the second solution will be \(y-2= y_1ln(x) \) so, \(y-2 = lnx + xlnx +\dfrac{lnx}{4}x^2+...... \) right?

OpenStudy (loser66):

\(y_2\) not y-2 hihihi

OpenStudy (anonymous):

Well, it should come out to whatever I had in the first post: \[y_{2}(x)= y_{1}(x)lnx - 2x - \frac{3}{4}x^{2}-\frac{11}{108}x^{3}+...\]

OpenStudy (loser66):

why?? is it not that we have formula to find the second solution is \(y_2 =y_1lnx\) ???

OpenStudy (loser66):

if we have double root from C.E?

OpenStudy (anonymous):

\[y_{2}(x) = y_{1}(x)lnx+x^{r_{1}}\sum_{n=1}^{\infty}b_{n}x^{n}\] Basically, Im unable to derive the correct coefficients in that series portion of the answer, the bn's

OpenStudy (loser66):

Wonder where is prof @SithAndGiggle ???

OpenStudy (anonymous):

Ive tried several different things to get those correct bn's, just not coming out correct.

OpenStudy (loser66):

oh, this is the second type of regular singular point. Since x =0 is not rg. sing. point, we find the value near that point. hihihi. However, your formula for the second solution is not correct. From my book it says \( y_2(x) = y_1(x)ln|x|+|x|^{r_1}\sum_{n=1}^\infty b_n(r_1)x^n\) where \(r_1=r_2~~(double~~root)\) So that the sum =0 when you have r =0. It turns to my solution \(y_2=y_1ln|x|\)

OpenStudy (anonymous):

Your formula is the same as mine, you just typed it more formally. My book has a sort of shortened version like I posted and the same thing you posted. Either way, I only know the solution my textbook has, which is what I listed above x_x So you're saying the answer in the back of my book is incorrect?

OpenStudy (amistre64):

this looks like a power series solution to me

OpenStudy (amistre64):

let: \[y=\sum_0a_nx^n\]\[y'=\sum_1a_n~nx^{n-1}\]\[y''=\sum_2a_n~n(n-1)x^{n-2}\] this gives us: \[x\sum_2a_n~n(n-1)x^{n-2} + \sum_1a_n~nx^{n-1} - \sum_0a_nx^n = 0\] even out the powers on x by shifting indexes: \[\sum_2a_n~n(n-1)x^{n-1} + \sum_1a_n~nx^{n-1} - \sum_1a_{n-1}x^{n-1} = 0\] now line up the starting summations by pulling out an n: \[a_1 -a_{0}+\sum_2a_n~n(n-1)x^{n-1} + \sum_2a_n~nx^{n-1} - \sum_2a_{n-1}x^{n-1} = 0\] and combine the summations by factoring out the x power \[a_1 -a_{0}+\sum_2~[a_n~n(n-1)+ a_n~n -a_{n-1}]~x^{n-1} = 0\] this is identically 0 when all the coefficients are 0, and we can devise a recurrsion equation now. \[a_n~n(n-1)+ a_n~n -a_{n-1}=0\] \[a_n~(n(n-1)+n) -a_{n-1}=0\] \[a_n~(n(n-1)+n) =a_{n-1}\] \[a_n =\frac{a_{n-1}}{n(n-1)+n}~:~n\ge2\]

OpenStudy (amistre64):

that denom looks like it can simplify n(n-1)+n n(n-1+1) = n^2

OpenStudy (anonymous):

x2y"+xy'-xy=0 r2+(1-1)r+0=0...so r2=0...r=0,0 then y1=sigma(an.x^n) y2=y1lnx+sigma(bnx^n)

OpenStudy (anonymous):

x2y"--->n(n-1)an xy'---->nan xy---->a(n-1)...n-1 is index so recursive equation is going to be...suppose a0=1 n(n-1)an+nan-a(n-1)=0----->an/a(n-1)=-1/n2------>an=(-1)^n/n!^2 then y1=sigma((-1)^n/n2.x^n)

OpenStudy (anonymous):

@amistre64 How do I get a second solution out of that, though? :(

OpenStudy (anonymous):

@mahmit2012 That gives me y1, which I already have been already able to get. How about y2? I don't know how to get the bn coefficients in that 2nd solution.

OpenStudy (amistre64):

a0 and a1 are your .... what are they called ..... constants for determining solutions. as is, you get a set of coefficients in terms of a0, and another in terms of a1

OpenStudy (amistre64):

or it may very well be that a0 is a constant, and the rest of the terms are in a1 .... havent worked it out too far yet